# Passive Beam Shaping via Binary-Coded Apertures

### v2

@@ VII-A4 Aperture-matched comparison
VII-A6 Energy-efficiency implication under surface-power overhead
@@
Fig. 13 illustrates the impact of surface-side operational power on energy efficiency at the target direction $\theta_{T}=-10^{\circ}$. Using the definitions in Section VI-B, we evaluate EE under a fixed radio-side power budget $P_{0}\triangleq P_{\mathrm{TX}}+P_{\mathrm{circ}}$ and vary only the surface-side term $P_{\mathrm{surf}}$. The achievable rate is obtained from the target-direction SNR using (32). The example parameters used to generate the curves are given in the figure caption (e.g., $B=2$ GHz and baseline $\mathrm{SNR}_{0}=-10$ dB). For each reflector design, the target-direction $\mathrm{SNR}(\theta_{T})$ is scaled according to its relative target gain, so that the resulting curves capture the joint effect of reflector-induced SNR improvement and added surface-side power.
@@
Figure 13: Energy efficiency $\mathrm{EE}=R/(P_{0}+P_{\mathrm{surf}})$ versus surface operational power $P_{\mathrm{surf}}$ at $\theta_{T}=-10^{\circ}$ (example setting: $P_{0}=1$ W, $B=2$ GHz, baseline $\mathrm{SNR}_{0}=-10$ dB). Fully passive designs have $P_{\mathrm{surf}}=0$ and thus yield constant $\mathrm{EE}$, while a powered RIS baseline decreases with $P_{\mathrm{surf}}$.
@@
VII-B Experimental Setup and OTA Results
VII-B1 Experimental setup and processing
OTA measurements are performed at 60 GHz using two Sivers EVK02001 transceiver kits on a constant-radius arc (0.9m) centered on the reflector. The transmitter is fixed at azimuth AoA $\theta_{I}$, while the receiver is swept in azimuth to measure the scattered response versus AoD $\theta_{T}$. Two reference scans are collected: mount-only (background) and all-ON (reference reflector). Background-subtracted power is computed as
@@ VII-B1 Experimental setup and processing
Figure 14: OTA azimuthal sweep measurement setup. The transmitter (TX) is fixed at the prescribed incidence angle, while the receiver (RX) is swept in azimuth to measure the scattered response.
VII-B2 OTA results and discussion
We evaluate both single-beam and multi-beam passive designs. The single-beam case uses $\theta_{I}=45^{\circ}$ with target $\theta_{T}=-10^{\circ}$, while the multi-beam case uses $\theta_{I}=30^{\circ}$ with targets $\theta_{T}\in\{-7.8^{\circ},-60^{\circ}\}$. In addition to the fabricated all-ON inkwell reflector, we include a smooth copper plate as a baseline reference to verify that the all-ON metallization provides comparable specular behavior and to enable a consistent conductor baseline for comparison.
@@ VII-B2 OTA results and discussion
For the multi-beam case ($\theta_{I}=30^{\circ}$), the ON/OFF and diffraction-order designs produce substantial enhancement at both target angles ($\theta_{T}=-7.8^{\circ}$ and $60^{\circ}$), confirming successful deterministic non-specular beam formation in a fully passive implementation. At the specular direction ($\theta_{T}=-30^{\circ}$), both designs exhibit reduced response relative to all-ON, which is expected since energy is redirected toward the non-specular targets. We note here that the measured magnitude of these specular losses and target gains do not always match the worst-case theoretical bounds relative to an ideal continuous-phase surface, however, the experimentally relevant comparison is against the passive all-ON baseline, for which the measured target enhancements are consistently large. This is becuase the theory in Section VII-A is developed using a 1-D azimuthal array-factor abstraction, whereas the fabricated reflector is a 2-D $35\times 35$ lattice and the measurements are azimuth cuts of a 2-D scattering pattern. Additional aperture extent in the orthogonal dimension, elevation-dependent weighting, and indoor multipath can affect absolute levels and partially explain deviations from idealized 1-D predictions. Nevertheless, the measured ranking and trends across all-ON, ON/OFF, and DGO designs remain consistent with theory as shown in Fig. 12.
VIII Conclusion

### v3

@@ VIII-A6 Wideband Sensitivity of Fixed Passive Designs
VIII-A7 Energy-efficiency implication under surface-power overhead
@@
Fig. 19 illustrates the impact of surface-side (reflector) operational power on energy efficiency at the target direction $\theta_{T}=-10^{\circ}$. Using the definitions in Section VII-B, we evaluate EE under a fixed radio-side power budget $P_{0}\triangleq P_{\mathrm{TX}}+P_{\mathrm{circ}}$ and vary only the surface-side term $P_{\mathrm{surf}}$. The achievable rate is obtained from the target-direction SNR using (37). The example parameters used to generate the curves are given in the figure caption. For each reflector design, the target-direction $\mathrm{SNR}(\theta_{T})$ is scaled according to its relative target gain, so that the resulting curves capture the joint effect of reflector-induced SNR improvement and added surface-side power.
@@
Figure 19: Energy efficiency $\mathrm{EE}=R/(P_{0}+P_{\mathrm{surf}})$ versus surface operational power $P_{\mathrm{surf}}$ at $\theta_{T}=-10^{\circ}$ for an example setting with $P_{0}=1$ W, $B=2$ GHz, and baseline $\mathrm{SNR}_{0}=-10$ dB. Fully passive designs have $P_{\mathrm{surf}}=0$ and therefore yield constant $\mathrm{EE}$, whereas a powered RIS baseline decreases with $P_{\mathrm{surf}}$. For the diffraction-order curve, unconstrained indicates that the selected period is applied while allowing the physical aperture to expand.
@@
VIII-B Experimental Setup and Results
VIII-B1 Experimental setup and processing
Over the air measurements are performed using a narrowband continuous-wave (CW) excitation at 60.48 GHz, which serves as the design frequency for the fabricated reflectors shown in Fig. 7. Two Sivers EVK02001 transceiver kits are positioned on a constant-radius arc (0.9 m) centered on the reflector. The transmitter is fixed at azimuth AoA $\theta_{I}$, while the receiver is swept in azimuth to measure the scattered response versus AoD $\theta_{T}$. Two reference scans are collected, mount-only (background) and all-ON (reference reflector). Background-subtracted power is computed as
@@ VIII-B1 Experimental setup and processing
Since the experimental measurements use a narrowband CW signal, they validate passive beam shaping at the design frequency rather than broadband communication performance.
Figure 20: Over the air azimuthal sweep measurement setup. The transmitter (TX) is fixed at the prescribed incidence angle, while the receiver (RX) is swept in azimuth to measure the scattered response.
VIII-B2 Experimental Results and Discussion
We experimentally evaluate the performance of the proposed passive reflector designs for both single-beam and multi-beam scenarios. The single-beam configuration considers an incidence angle of $\theta_{I}=45^{\circ}$ with a target direction $\theta_{T}=-10^{\circ}$, while the multi-beam case uses $\theta_{I}=30^{\circ}$ with target directions $\theta_{T}\in\{-7.8^{\circ},-60^{\circ}\}$. In addition to the fabricated all-ON inkwell reflector, a smooth copper plate is included as a reference baseline to verify that the all-ON metallization achieves comparable specular reflection and to establish a consistent conductor benchmark for subsequent comparisons.
@@ VIII-B2 Experimental Results and Discussion
For the multi-beam case ($\theta_{I}=30^{\circ}$), the ON/OFF and diffraction-order designs produce substantial enhancement at both target angles ($\theta_{T}=-7.8^{\circ}$ and $60^{\circ}$), confirming successful deterministic non-specular beam formation in a fully passive implementation. At the specular direction ($\theta_{T}=-30^{\circ}$), both designs exhibit reduced response relative to all-ON, which is expected since energy is redirected toward the non-specular targets. We note here that the measured magnitude of these specular losses and target gains do not always match the worst-case theoretical bounds relative to an ideal continuous-phase surface, however, the experimentally relevant comparison is against the passive all-ON baseline, for which the measured target enhancements are consistently large. This is becuase the theory in Section VIII-A is developed using a 1-D azimuthal array-factor abstraction, whereas the fabricated reflector is a 2-D $35\times 35$ lattice and the measurements are azimuth cuts of a 2-D scattering pattern. Additional aperture extent in the orthogonal dimension, elevation-dependent weighting, and indoor multipath can affect absolute levels and partially explain deviations from idealized 1-D predictions. Nevertheless, the measured ranking and trends across all-ON, ON/OFF, and diffration order designs remain consistent with theory.
TABLE III: Measured response loss versus frequency for representative ON/OFF and diffraction-order reflectors, normalized to the 60.48 GHz design point, for $\theta_{I}=45^{\circ}$ and $\theta_{T}=-10^{\circ}$.
Frequency (GHz) ON/OFF loss (dB) Diffraction-order loss (dB)
58.32 -1.2 -3.3
60.48 0.0 0.0
62.64 -3.0 -5.4
64.80 -8.0 -11.5
66.90 -14.2 -16.3
69.10 -13.4 -20.7
To experimentally assess frequency sensitivity beyond the nominal 60.48 GHz design point, additional narrowband CW measurements were conducted at 58.32, 60.48, 62.64, 64.80, 66.90, and 69.10 GHz for representative ON/OFF and diffraction-order reflectors in the beam-shaping configuration $\theta_{I}=45^{\circ}$ and $\theta_{T}=-10^{\circ}$. The measured responses were normalized to the 60.48 GHz result, and the corresponding losses are summarized in Table III. The results show that the target-direction response at $\theta_{T}=-10^{\circ}$ degrades for both reflector types as the operating frequency shifts away from the design frequency. Across all measured frequencies, the diffraction-order reflector consistently exhibits larger loss than the ON/OFF design. For example, at 64.80 GHz the ON/OFF reflector experiences a loss of 8.0-8.0 dB compared to 11.5-11.5 dB for the diffraction-order design, while at 69.10 GHz the losses increase to 13.4-13.4 dB and 20.7-20.7 dB, respectively. This behavior is in close agreement with the analytical trends in Fig. 17, confirming that the ON/OFF design is comparatively more robust to frequency variation. This degradation can be understood in terms of the wavelength dependence of the underlying beamforming mechanisms. Diffraction-order steering relies explicitly on the ratio between the fixed spatial period and the wavelength (see (31)), such that changes in frequency shift the effective diffraction condition and reduce gain in the intended direction. In contrast, the ON/OFF design relies on dense spatial modulation of the aperture and is therefore less sensitive to moderate changes in wavelength, although it still experiences performance degradation away from the design point.
IX Conclusion